What is the electrostatic force of attraction between a -6.0 x 10^-6 C charge and a 4.0 x 10^-6 C if they are separated by a distance of 3 m ?


1
Expert's answer
2021-10-12T12:32:14-0400

"F=k\frac{q_1q_2}{x^2}=9\cdot10^9\cdot\frac{|6\cdot10^{-6}|\cdot|4\cdot10^{-6}|}{3^2}=0.024\ (N)"